16) Todd has a margin account with $17,400 in available cash. The initial margin is 70 percent and the maintenance margin is 30 percent. What is the maximum number of shares he can purchase if the price per share is $44?

17) Aaron purchased 300 shares of a technology stock for $16.80 a share. The initial margin requirement on this stock is 85 percent and the maintenance margin is 60 percent. What is the lowest the stock price can go before he receives a margin call?

18) Five months ago, you purchased 400 shares of stock on margin. The initial margin requirement on your account is 60 percent and the maintenance margin is 30 percent. The call money rate is 4.8 percent, plus 1.85 percent premium. The purchase price was $16 a share. Today, you sold these shares for $18.00 each. What is your holding period rate of return
Answer #1

1.
=17400/(70%*44)=564.9350649

2.
(300*P-(300*16.80*15%))/(300*P)<=60%
=>P<=6.296063

3.
=(400*18-400*16-400*16*(1-60%)*(4.8%+1.85%))/(400*16*60%)=16.400%